OpenStudy (anonymous):

1. An object falls from rest on a high tower and takes 5.0 s to hit the ground. Calculate the object’s position from the top of the tower at 1.0 s intervals. Make a position-time graph for the object’s motion. In your response, show what you are given, the equation that you used, any algebra required, a table of data, and your graph. g = 9.8 m/s2

OpenStudy (anonymous):

can anybody help?

OpenStudy (anonymous):

Equation of the falling motion: y = -gt^2/2, from the top of the building With g = 9.8 -> y = 4.9*t^2 At time t =1, the object is far from the top: y(1) = 4,9 m At time t = 2, y(2) = 4.9(*4) = 19.6 m At time t = 3: y(3) = 9*4.9 = 44.1 m At time t = 4: y(4) = 16*(4.5) = 78.4 At time t = 5: y(5) = 25*(4.9) = 122.5 The object reached the soil in 5 sec, then the height of the building is 122.5 m.